# Trading Devil Final: Backdoor attack via Stock market and Bayesian Optimization

### v1

@@ III Adversarial Machine Learning in Finance via Bayesian Approach
“MarketBackFinal 2.0” is a technique that implements a poisoning attack with a ”clean-label backdoor“ [58],[59],[60],[61],[62],[63]. Contains methods such as ‘stochastic investments models‘, rough volatility algorithm 3, optimal transport algorithm 5 (calculating the deterministic movement (transport component) based on the state and time), and dynamic hedging algorithm 6(which takes as calculate Call Option Price algorithm4 and calculate Greeks algorithm 7), Bayesian Optimization Process algorithm 8, Diffusion Bayesian Optimization algorithm 1 (which takes as setting up hierarchical priors algorithm 2 ) to apply the attack to the audio data, Bayesian style is implemented using a “prior” (Bayesian optimization over the given objective function) and the BO 151515bayesian-optimization framework with drifts functions including Diffusion Bayesian Optimization. Thanks to this technique, process volatility effects in the drift function (incorporating the transport component into the drift functions) are used for sampling to obtain and define the prior distribution, and a diffusion technique is then applied, which implements a diffusion-based sampling technique to generate a sequence of samples as a function of certain parameters and a noise distribution. The Bayesian method integrates the drift function into the Bayesian model in the “MarketBackFinal 2.0” method while using a NUTS method for sampling (efficient sampling with adaptive step size Hamiltonian Monte Carlo). Given a time step $T$ and a set of parameters $\alpha,\beta,\sigma,\theta$, the method generates a new data point $x_{T}$ based on the current state $x_{T-1}$ and the noise distribution $sin(x)$. The results are available on ART.1.18; link: https://github.com/Trusted-AI/adversarial-robustness-toolbox/pull/2467.
@@ Conclusions
The weaknesses of transformer-based generative artificial intelligence models are the main subject of this work, which also showcases a new financial simulation tool. A clean-label backdoor and poisoning attack specifically designed for financial modeling using various inversion models via diffusion drift optimized by Bayesian[74] simulation is known as “MarketBackFinal 2.0,” and it is one of the attack methods developed in this article. Function simulations utilize hierarchical hypothesis parameters. With the primary goal of using them with financial data, this article concentrates on the development of new financial simulation tools. But in order to ensure that the approach method operates as intended, temporal acoustic data obtained through a backdoor assault is used for validation. Audio backdoor attacks based on Bayesian transformations (using a drift function via stochastic investment model effects for sampling [75]) based on a diffusion model approach (which adds noise). The study results help to understand the potential but also the risks and vulnerabilities to which pre-trained advanced DNN models are exposed via malicious audio manipulation to ensure the security and reliability of automatic speech recognition audio models.

### v2

@@ III Adversarial Machine Learning in Finance via Bayesian Approach
“MarketBackFinal 2.0” is a technique that implements a poisoning attack with a ”clean-label backdoor“ [58],[59],[60],[61],[62],[63],[64]. Contains methods such as ‘stochastic investments models‘, rough volatility algorithm 3, optimal transport algorithm 5 (calculating the deterministic movement (transport component) based on the state and time), and dynamic hedging algorithm 6(which takes as calculate Call Option Price algorithm4 and calculate Greeks algorithm 7), Bayesian Optimization Process algorithm 8, Diffusion Bayesian Optimization algorithm 1 (which takes as setting up hierarchical priors algorithm 2 ) to apply the attack to the audio data, Bayesian style is implemented using a “prior” (Bayesian optimization over the given objective function) and the BO 151515bayesian-optimization framework with drifts functions including Diffusion Bayesian Optimization. Thanks to this technique, process volatility effects in the drift function (incorporating the transport component into the drift functions) are used for sampling to obtain and define the prior distribution, and a diffusion technique is then applied, which implements a diffusion-based sampling technique to generate a sequence of samples as a function of certain parameters and a noise distribution. The Bayesian method integrates the drift function into the Bayesian model in the “MarketBackFinal 2.0” method while using a NUTS method for sampling (efficient sampling with adaptive step size Hamiltonian Monte Carlo). Given a time step $T$ and a set of parameters $\alpha,\beta,\sigma,\theta$, the method generates a new data point $x_{T}$ based on the current state $x_{T-1}$ and the noise distribution $sin(x)$. The results are available on ART.1.18; link: https://github.com/Trusted-AI/adversarial-robustness-toolbox/pull/2467.
@@ Conclusions
The weaknesses of transformer-based generative artificial [75] intelligence models are the main subject of this work, which also showcases a new financial simulation tool. A clean-label backdoor and poisoning attack specifically designed for financial modeling using various inversion models via diffusion drift optimized by Bayesian[76],[77] simulation is known as “MarketBackFinal 2.0,” and it is one of the attack methods developed (Figure 15) in this article. Function simulations utilize hierarchical hypothesis parameters. With the primary goal of using them with financial data, this article concentrates on the development of new financial simulation tools. But in order to ensure that the approach method operates as intended, temporal acoustic data obtained through a backdoor assault is used for validation. Audio backdoor attacks [78], [79],[80], [81], [82],based on Bayesian transformations (using a drift function via stochastic investment model effects for sampling [83]) based on a diffusion model approach [84] (which adds noise). The study results help to understand the potential but also the risks and vulnerabilities to which pre-trained advanced DNN models are exposed via malicious audio manipulation to ensure the security and reliability of automatic speech recognition audio models. MarketBackFinal 2.0 exposes vulnerabilities arising from the diffusion models developed in [85], [86], [87], [88], [89],[6],[90],[91].
Figure 15: 3D visualization surface
@@
Self-financing: The value $V_{t}$ of a portfolio containing $\psi_{t}$ units of bonds and $\phi_{t}=\left(\phi_{t,1},\ldots,\phi_{t,m}\right)^{\top}$ units of various stocks at a given time $t$ is given by
$V_{t}=B_{t}\psi_{t}+\sum_{i=1}^{m}\phi_{t,i}S_{t,i}=B_{t}\psi_{t}+\phi_{t}^{\top}\mathbf{S}_{t}.$
The particular value of the portfolio at time $t$ typically depends on the units of assets $\left(\psi_{t},\phi_{t}\right)$ held just before time $t$. The stochastic process $\left\{\left(\psi_{t},\phi_{t}\right),t\geqslant 0\right\}$ can thus be interpreted as a trading strategy.
$\left\{\mathcal{F}_{t},t\geqslant 0\right\}$.
$\mathrm{d}V_{t}=\left(B_{t}\mathrm{\leavevmode\nobreak\ d}\psi_{t}+\mathbf{S}_{t}^{\top}\mathrm{d}\phi_{t}\right)+\psi_{t}\mathrm{\leavevmode\nobreak\ d}B_{t}+\phi_{t}^{\top}\mathrm{d}\mathbf{S}_{t}$
The portfolio is called self-financing if
$B_{t}\mathrm{\leavevmode\nobreak\ d}\psi_{t}+\mathbf{S}_{t}^{\top}\mathrm{d}\phi_{t}=0$
in which case
$V_{T}=V_{0}+\int_{0}^{T}\psi_{t}\mathrm{\leavevmode\nobreak\ d}B_{\ell}+\int_{0}^{T}\phi_{t}^{\top}\mathrm{d}\mathbf{S}_{t}.$
Arbitrage-free market: A market model is called arbitrage-free on $[0,T]$ if, in essence, for every self-financing trading strategy $\left\{\left(\psi_{t},\phi_{t}\right),0\leqslant t\leqslant T\right\}$,
$V_{0}=0\Rightarrow\mathbb{P}\left(V_{T}>0\right)=0\text{. }$
$\left\{\widetilde{\mathbf{S}}_{t},0\leqslant t\leqslant T\right\}$, with $\widetilde{\mathbf{S}}_{t}=B_{t}^{-1}\mathbf{S}_{t}$, is an $\left(\mathcal{F}_{t},\mathbb{Q}\right)$-martingale.
$\sigma_{t}\mathbf{u}_{t}=\mu_{t}-r_{t}\mathbf{S}_{t}$
has a solution $\mathbf{u}_{t},t\in[0,T]$ for which $\mathbb{E}\exp\left(\frac{1}{2}\int_{0}^{T}\mathbf{u}_{t}^{\top}\mathbf{u}_{t}\mathrm{\leavevmode\nobreak\ d}t\right)<\infty$
$\mathrm{d}\widetilde{\mathbf{S}}_{t}$ $=B_{t}^{-1}\left(\boldsymbol{\mu}_{t}-r_{t}\mathbf{S}_{t}\right)\mathrm{d}t+B_{t}^{-1}\sigma_{t}\mathrm{\leavevmode\nobreak\ d}\mathbf{W}_{t}$
$=B_{t}^{-1}\sigma_{t}\left(\mathbf{u}_{t}\mathrm{\leavevmode\nobreak\ d}t+\mathrm{d}\mathbf{W}_{t}\right)$
$=B_{t}^{-1}\sigma_{t}\mathrm{\leavevmode\nobreak\ d}\mathbf{Z}_{t}$
where $\mathrm{d}\mathbf{Z}_{t}=\mathbf{u}_{t}\mathrm{\leavevmode\nobreak\ d}t+\mathrm{d}\mathbf{W}_{t}$ defines an Itô process. Let
$M_{t}=\exp\left(\int_{0}^{t}\mathbf{u}_{s}^{\top}\mathrm{d}\mathbf{W}_{s}-\frac{1}{2}\int_{0}^{t}\mathbf{u}_{s}^{\top}\mathbf{u}_{s}\mathrm{\leavevmode\nobreak\ d}s\right).$
Then, by Girsanov’s theorem [95], $\left\{M_{t},t\geqslant 0\right\}$ is an $\left(\mathcal{F}_{t},\mathbb{P}\right)$-martingale, and under the new measure $\mathbb{Q}(A)\stackrel{{\scriptstyle\text{ def }}}{{=}}\mathbb{E}\left[M_{T}\mathrm{I}_{A}\right]$ for all $A\in\mathcal{F}_{T}$, the process $\left\{\mathbf{Z}_{t},0\leqslant\right.$ $t\leqslant T\}$ is a Wiener process. It follows that under $\mathbb{Q}$ the process $\left\{\widetilde{\mathbf{S}}_{t},0\leqslant t\leqslant T\right\}$ is a martingale with respect to $\left\{\mathcal{F}_{t},0\leqslant t\leqslant T\right\}$, which had to be shown. $\mathbb{Q}$ is called the risk-neutral measure.
Let $\widetilde{V}_{t}=B_{t}^{-1}V_{t}=\psi_{t}+\boldsymbol{\phi}_{t}^{\top}B_{t}^{-1}\mathbf{S}_{t}=\psi_{t}+\boldsymbol{\phi}_{t}^{\top}\widetilde{\mathbf{S}}_{t}$ be the discounted value of the portfolio of stocks and bonds.
$\mathrm{d}\widetilde{V}_{t}=\boldsymbol{\phi}_{t}^{\top}\mathrm{d}\widetilde{\mathbf{S}}_{t}=B_{t}^{-1}\boldsymbol{\phi}_{t}^{\top}\boldsymbol{\sigma}_{t}\mathrm{\leavevmode\nobreak\ d}\mathbf{Z}_{t}$
$\mathbb{Q}$ the process $\left\{\tilde{V}_{t},0\leqslant t\leqslant T\right\}$ is a martingale with respect to $\left\{\mathcal{F}_{t},0\leqslant t\leqslant T\right\}$. As a consequence,
$V_{t}=B_{t}\mathbb{E}_{\mathbb{Q}}\left[B_{T}^{-1}V_{T}\mid\mathcal{F}_{t}\right],\quad t\leqslant T.$
The payoff at maturity for a call option is $\left(S_{T}-K\right)^{+}=\max\left\{S_{T}-K,0\right\}$ and for a put option is $\left(K-S_{T}\right)^{+}$. Suppose for concreteness that we are dealing with a call option and denote its value at time $t\leqslant T$ by $C_{t}=C\left(S_{t},K,r_{t},T-t\right)$, where $T-t$ is the time until maturity.
$C_{t}=V_{t}=B_{t}\mathbb{E}_{\mathbb{Q}}\left[B_{T}^{-1}C_{T}\mid\mathcal{F}_{t}\right],\quad t\leqslant T,$
where
$C_{T}=V_{T}=V_{0}+\int_{0}^{T}\psi_{t}\mathrm{\leavevmode\nobreak\ d}B_{t}+\int_{0}^{T}\boldsymbol{\phi}_{t}^{\top}\mathrm{d}\mathbf{S}_{t}.$
Option prices provide information about state variables and parameters via the pricing equation. An option price is given by
$C\left(S_{t},V_{t},\theta\right)=e^{-r(T-t)}\mathbb{E}_{t}^{\mathbb{Q}}\left[\max\left(S_{t}-K,0\right)\mid S_{t},X_{t},\Theta\right]$
where the expectation is taken under $\mathbb{Q}$ the risk-neutral probability measure. here we assume that the risk-free rate $r$ is constant. This simplifies by letting $A=\left\{S_{T}\geq K\right\}$ denote the event that the stock ends in the money. Then the option price is given by :
$e^{-rt}\mathbb{E}_{t}^{\mathbb{Q}}\left(\max\left(S_{T}-K,0\right)\right)$ $=e^{-rt}\mathbb{E}_{t}^{\mathbb{Q}}\left(S_{T}\mathbb{I}_{A}\right)-e^{-rt}K\mathbb{E}\left(\mathbb{I}_{A}\right)$
$=e^{-rt}\mathbb{E}_{t}^{\mathbb{Q}}\left(S_{T}\mathbb{I}_{A}\right)-e^{-rt}K\mathbb{P}(A).$
The logarithmic asset price and volatility follow an affine process with $Y_{t}=\left(\log S_{t},V_{t}\right)\in\mathfrak{R}\times\mathfrak{R}^{+}$satisfying
$dY_{t}=\left(r-\frac{1}{2}V_{t}\right)dt+\sqrt{V_{t}}dB_{1,t}$
$dV_{t}=\kappa\left(\theta-V_{t}\right)dt+\sigma_{v}\sqrt{V_{t}}dZ_{t}$
where $Z_{t}=\rho B_{1,t}+\sqrt{1-\rho^{2}}B_{2,t}$. The correlation, $\rho$, or so-called leverage effect is important to explain the empirical fact that volatility increases faster as equity prices drop. The parameters $(\kappa,\theta)$ govern the speed of mean reversion and the long-run mean of volatility and $\sigma_{v}$ measures the volatility of volatility. Under the risk-neutral measure they become $\left(\kappa/\kappa+\lambda_{v}\right)\theta$ and $\kappa+\lambda_{v}$ where $\lambda_{v}$ is the market price of volatility risk.
From affine process theory, the discounted transform
$\psi(u)=e^{-rt}\mathbb{E}\left(e^{uY_{T}}\mid Y_{t}\right)=e^{\alpha(t,u)+uY_{t}+\beta(t,u)V_{t}}$
where $\alpha,\beta$ satisfy Riccati equations.